# A reduced-order framework for temperature estimation in food freezing from optimally located sensors, including turbulent conjugate flow scenarios

### v1

@@ Abstract
This article proposes a framework for estimating temperature fields in food-freezing applications that significantly reduces computational load while ensuring accurate temperature monitoring, representing a promising technological tool for optimizing and controlling food engineering processes. The strategy is based on (i) a mathematical model of a convection-dominated problem coupling thermal convection and turbulence and (ii) a least-squares approach for solving the inverse data assimilation problem, regularized by projecting the governing dynamics onto a reduced-order model (ROM). The unsteady freezing process considers an idealized salmon slice in a freezer cabinet, modeled with temperature-dependent thermophysical properties. The forward problem is approximated using a third-order WENO finite volume solver, including an optimized second-order backward scheme for time discretization. We employ our data assimilation framework to reconstruct the temperature field from a limited number of sensor data and to estimate temperature distributions within frozen food. Sensor placement is optimized using a new greedy algorithm, relying on maximizing the observability of the reduced-order dynamics for a fixed set of sensors. The proposed approach allows efficient extrapolation from external sensor measurements to the internal temperature of the food, which is crucial for maintaining food quality.
@@
Food Freezing , Data assimilation , Greedy Algorithm , Reduced Order Modelling , Phase Change , Optimal Sensor Placement
@@ 1 Introduction
Freezing processes, crucial for transporting and preserving food [1, 2], have important economic implications and rely on large amounts of energy required to satisfy safe freezing conditions. Food freezing involves a highly nonlinear liquid-solid phase-change that produces significant changes in the processed food and strongly depends on the operational conditions imposed, for example, in a freezing tunnel. Liquid-solid phase change implies a boundary evolution scenario in which the phases interact, natural convection occurs, the thermophysical properties of the food change strongly, and turbulence develops in the airflow that freezes the food [3, 4]. In addition, food humidity, texture, color, and flavor depend on the evolution of the freezing rate [5, 6, 7]. All these factors influence food quality and the possibility of estimating in advance how the food temperature changes during freezing is of both scientific and technological interest.
@@
Computational fluid dynamics has become a standard tool to support the understanding of the physics of fluids and of the heat-mass transfer that occurs in food processing. For example, computer simulations have been successfully used for evaluating convective coefficients [8, 9, 10], for estimating freezing times [11, 12, 13], for assessing food positioning for better freezing [3], and for testing the effectivity of different operational conditions [14], thereby improving the overall thermal processes. Numerical methods used for these simulations have been based on finite difference [15, 16], finite element [17, 18, 19] or finite volume methods [20, 21, 22]. The advantages and disadvantages of these numerical methodologies have been recently discussed in [23, 24].
@@
Inverse problems in the food industry have gained attention in recent years, primarily focusing on parameter estimation techniques. Recent studies, such as those in [25], emphasize the use of inverse strategies, particularly in identifying material properties and diffusivity parameters in food products. For example, research in [26] highlights the estimation of moisture diffusivity by inverting a finite element model applied to various solid foods such as salami, biscuits, and flatbread by minimizing the distance between the numerical model and experimental results through the Levenberg–Marquardt algorithm. Just as ensuring the properties of heat-treated foods is of scientific and technological interest, efficient energy consumption in the processes is also an aspect that may be optimized using similar techniques [27]. In [28], the authors highlight that using the bootstrap technique can estimate both specific heat and thermal conductivity in a freezing range, comparing the obtained results with experimental data.
In this work, we address the estimation of the temperature fields in a freezing salmon slice from measurement data, focusing on the case of partial and low-resolution observations. This type of research belongs to the broader field of data assimilation, where available data and mathematical models are used both to estimate unknown parameters of a PDE (parameter estimation) or to correct model results (state estimation) [29, 30, 31]. In the latter case, as the problem considered in this work, the goal is to estimate time- and space-resolved fields over the whole computational domain or on a subset of interest.
@@
For the forward modeling, we use a third-order weighted essentially non-oscillatory finite volume method (WENO3) combined with an optimized second-order backward differentiation formulation (BDF2-opt) with an implicit scheme to handle the temporal evolution. This approach ensures both accuracy and stability. Efficient implementations of these methods and high-performance computing techniques allow accurate and fast approximation of thermally-coupled flows in complex geometries. However, the need to systematically repeat these simulations, a key stage for the inverse problem, calls for the usage of Reduced-Order Modeling (ROM) strategies to reliably decrease the dimensionality of the problem, preserving its fundamental characteristics, using techniques such as the reduced-basis method [32], or proper orthogonal decompositions [33] (see e.g. [34, 35], for a review on ROM strategies).
@@
As an inverse technique, we employ a regularized least-squares approach based on the so-called Parametrized Background Data Weak (PBDW) method, see, e.g., [36, 37, 38, 39]. In this approach, the optimization problem to be solved for the state estimation is based on the projection of the physics on a reduced-dimensional space that must reproduce the observed data in a weak form. Application of the PBDW in biomedical problems have been recently presented in [40, 41, 42, 43]. In this work, we use a methodology inspired by the PBDW to assimilate the temperature fields in the liquid-solid phase-change scenario, including turbulent flow conditions. In particular, we employ the ROM-based component of PBDW, which is equivalent to the full PBDW approach with infinite model regularization, as explained in [38], assuming no bias on the physical model. Additionally, we combine the state estimation with a Greedy algorithm for the optimal placement of temperature sensors. Greedy algorithms have been used for experimental design in the past years. In [44], a Bayesian design is used for seismic full waveform inversion. In [45], a greedy algorithm was employed using a minimal variance error for parameter estimation. In [46], optimal sensors are computed targeting directly a mean squared error between the model and measurements. Greedy methods have been combined with the PBDW in, e.g., [47, 48, 38]. In contrast with these works, our proposal solely relies on available a priori results for the regularized least-squares reconstruction, prioritizing sensor locations that maximize the observability of the reduced-order model.
@@
In summary, this article proposes a temperature estimation framework from optimally placed synthetic data and a regularized least-squares estimator through a physics-informed projection, using turbulent simulations of the airflow surrounding the freezing food. The temperature estimation algorithm is proven valid for extrapolating fields in non-observable regions, keeping good reconstruction errors due to a properly designed sensor placement strategy. As a further application of the fully reconstructed field, we also show the capabilities of the algorithm to predict temperature-related quantities, such as Nusselt numbers, freezing time, and convective coefficients, aspects of high interest in food engineering. This proof-of-concept article will specifically use synthetically generated measurements from a thermal camera.
@@ 2.1 Model setting
We model the freezing process of food inside a convective freezer that circulates cold air between the cabinet and the refrigeration unit through inlet and outlet ducts. The conjugate dynamics of the process result from the simultaneous interplay of forced heat convection in the turbulent airflow and the nonlinear heat transfer coupled with the liquid-to-solid phase change process within the food slab. This setup is illustrated in Figure 1.
The freezer is assumed to have a cabinet of width $W=0.35$ m, height $H=0.4$ m, and depth $D=0.45$ m. Cold air coming from the refrigeration unit is forced into the cabinet by a fan at a constant velocity $U_{in}$ and temperature $T_{C}$ through an inlet duct of height $H_{in}=0.12$ m, and it leaves the cabinet through an outlet duct of height $H_{out}=0.06$ m. An idealized hexahedral food slab of width $W_{f}=0.08$ m, height $H_{f}=0.03$ m, and depth $D_{f}=0.08$ m, is placed on a glass shelf of width $W_{s}=0.25$ m and thickness $d_{s}=0.002$ m, located at the mid-height of the cabinet. Also, we assume that the walls of the cabinet exchange heat with the environment by conduction through the insulated layers and by natural convection with the surrounding air. The rear wall that separates the cabinet from the refrigeration unit is considered adiabatic.
Figure 1: Left: Sketch of the physical domain for the food freezing process. Right: Details of the two-dimensional section considered in this study.
In the present work, we focus on using an inverse pipeline to recover temperature fields from low-quality measurements that require building a dataset with multiple direct simulations. Therefore, we only consider a two-dimensional setup at the mid-section of the freezer (Figure 1, right). This two-dimensional model has been proven to represent with sufficient accuracy the food freezing process by turbulent forced and natural convection of airflow, see, e.g., the specialized references [10, 49]. In what follows, we will denote with $\Omega_{f}\subset\mathbb{R}^{2}$ and $\Omega_{s}\subset\mathbb{R}^{2}$ the fluid-flow and food domains, respectively, and $[0,t_{f}]$ is the physical time interval considered for the simulation.
@@ 2.2 Airflow dynamics
$\rho\partial_{t}T+\rho\boldsymbol{u}\cdot\nabla T-\nabla\cdot[(\lambda/C_{p}+\sigma_{T}\mu_{t})\nabla T]=0\qquad\text{in}\ \Omega_{f}\times(0,t_{f}],$
In Eq. (1) $\rho$, $\lambda$, $C_{p}$, $\mu$, and $\beta_{ref}$ denote the density, thermal conductivity, specific heat, and viscosity of the air, respectively. These physical properties are assumed as constant at a reference temperature of $T_{ref}=0$$\degree$°C. The source term in the momentum equation,
@@
includes the buoyancy and the turbulent source terms, where $\beta_{ref}$ is the air’s thermal expansion coefficient and $k$ is the turbulent kinetic energy. The additional turbulent viscosity $\mu_{t}$ introduced by the URANS method is described by the $k$-$\omega$ SST model as a function of the turbulent kinetic energy and specific turbulence dissipation rate ($\omega$) as
$\mu_{t}=\rho\,\frac{a^{*}k}{\max(a_{1}\omega,SF_{2})}\,,$ (2)
while two additional scalar partial differential equations are introduced to close the model:
$\rho\partial_{t}k+\rho\boldsymbol{u}\cdot\nabla k-\nabla\cdot(\mu+\sigma_{k}\mu_{t})\nabla k=\tilde{P}_{k}-\beta^{*}\rho k\omega\qquad\text{in}\ \Omega_{f}\in(0,t_{f}]$ (3)
$\rho\partial_{t}\omega+\rho\boldsymbol{u}\cdot\nabla\omega-\nabla\cdot(\mu+\sigma_{\omega}\mu_{t})\nabla\omega=\frac{\rho\gamma}{\mu_{t}}\tilde{P}_{k}-\rho\beta\omega^{2}+C_{k\omega}\qquad\text{in}\ \Omega_{f}\in(0,t_{f}].$
@@ 2.3 Food freezing process
The heat transfer and the liquid-solid phase change process in the food slab are described by a non-linear energy equation, assuming an incompressible and homogeneous material without considering mass transfer, such that:
$\partial_{t}(\rho_{s}C_{s}T_{s})-\nabla\cdot(\lambda_{s}\nabla T_{s})=0\qquad\text{in}\ \Omega_{s}\in(0,t_{f}].$ (4)
The internal convection of the liquid water contained in the muscle fiber and connective tissue of salmon can be neglected at the macro-level [20]. To account for the effect of latent heat during the phase change process in the salmon slab, the thermal properties (the density $\rho_{s}$, the specific heat capacity $C_{s}$, and the thermal conductivity $\lambda_{s}$) are considered as temperature-dependent, according to the effective heat capacity method [51, 52]. The temperature variability of the properties is depicted in Figure 2, assuming third-order fitting with (piecewise-constant) temperature-dependent coefficients:
@@ 2.4 Boundary, initial, and interface conditions
The inverse pipeline of the present work requires a dataset of snapshots from a set of time-dependent direct simulations with a variable space of parameters. Therefore, the magnitudes of the main boundary conditions are varied within a fixed range: $U_{in}\in[0.15,0.25]$ m-1, $T_{C}\in[-18,-26]\ ^{\circ}\text{C}$, $T_{ext}\in[18,26]\ ^{\circ}\text{C}$, and $h_{ext}\in[0.4,1.2]\ \text{W}\text{m}^{-2}\text{K}^{-1}$. The detailed procedure to construct this five-dimensional set is explained in Section 5.

### v2

# A fast food-freezing temperature estimation framework using optimally located sensors

@@ Abstract
This article presents and assesses a framework for estimating temperature fields in real time for food-freezing applications, significantly reducing computational load while ensuring accurate temperature monitoring, which represents a promising technological tool for optimizing and controlling food engineering processes. The strategy is based on (i) a mathematical model of a convection-dominated problem coupling thermal convection and turbulence, and (ii) a least-squares approach for solving the inverse data assimilation problem, regularized by projecting the governing dynamics onto a reduced-order model (ROM). The unsteady freezing process considers a salmon slice in a freezer cabinet, modeled with temperature-dependent thermophysical properties. The forward problem is approximated using a third-order WENO finite volume solver, including an optimized second-order backward scheme for time discretization. We employ our data assimilation framework to reconstruct the temperature field based on a limited number of sensors and to estimate temperature distributions within frozen food. Sensor placement is optimized using a novel greedy algorithm, which maximizes the observability of the reduced-order dynamics for a fixed set of sensors. The proposed approach allows efficient extrapolation from external sensor measurements to the internal temperature of the food under realistic turbulent flow conditions, which is crucial for maintaining food quality.
@@
Food Freezing , Phase Change , Inverse Problemas , Reduced order Modelling , Greedy Algorithm , Optimal Sensor Placement
@@ 1 Introduction
Freezing processes are crucial for food transport and preservation [1, 2] and have important economic implications, relying on large amounts of energy to satisfy safe freezing conditions. Food freezing involves a highly nonlinear liquid-solid phase change that produces significant variation in processed food and strongly depends on the operational conditions imposed, for example, in a freezing tunnel. Liquid-solid phase change implies a boundary evolution scenario in which the phases interact, natural convection occurs, the thermophysical properties of the food change strongly, and turbulence develops in the airflow that freezes the food [3, 4, 5]. In addition, food humidity, texture, color and flavor depend on the evolution of the freezing rate [6, 7, 8]. All these factors influence food quality, and the possibility of estimating in advance how the food temperature changes during freezing is of both scientific and technological interest.
@@
Computational fluid dynamics has become a standard tool to support the understanding of the physics of fluids and of the heat-mass transfer that occurs in food processing. For example, computer simulations have been successfully used to evaluate convective coefficients [9, 10, 11], to estimate freezing times [12, 13, 14], and to test the effectivity of different operational conditions [15], thus improving overall thermal processes. Numerical methods used for these simulations have been largely based on finite difference [16, 17], finite element [18, 19, 20] or finite volume methods [21, 22]. The strengths and weaknesses of each of these numerical methodologies have been recently discussed in [23, 24, 25].
@@
Inverse problems in the food industry have gained attention in recent years, primarily focusing on parameter estimation techniques. Recent studies, such as those in [26], emphasize the use of inverse strategies, particularly in identifying material properties and diffusivity parameters in food products. For example, research in [27] highlights the estimation of moisture diffusivity by inverting a finite element model applied to various solid foods through the Levenberg–-Marquardt algorithm. Just as ensuring the properties of heat-treated foods is of scientific and technological interest, efficient energy consumption in the processes is also an aspect that can be optimized using similar techniques [28]. In [29], comparison with experimental data showed that the bootstrap technique can estimate both specific heat and thermal conductivity in a freezing range.
In this work, we address the estimation of the temperature fields in a freezing salmon slice from measured data, focusing on the case of partial and low-resolution observations. This type of research belongs to the broader field of data assimilation, where available data and mathematical models are used both to estimate unknown parameters of a PDE (parameter estimation) [30] or to fully discover model fields (state estimation) [31, 32, 33]. A variety of data-driven methods have emerged recently in the literature, powered up by the gathering of large data-sets. Nevertheless, this kind of approach, though undoubtedly capable, may suffer drawbacks in scenarios where there is much less access to repetitive measurements, or whenever physical interpretability is a must [34]. This calls for alternatives that find a compromise between the purely data-driven approach and the physically informed approach, thus exploiting some well-known governing laws of the underlying phenomena. [35, 36, 37, 38, 39]. In the latter case, as in the problem considered in this work, the goal is to estimate time- and space-resolved fields over the whole computational domain or on a subset of interest, starting from experimental measurements that are typically noisy. This usually leads to ill-conditioned problems, and the literature propose several strategies to convey such estimations, such as maximally weighted approaches [40] or classical Kalman filters [41].
@@
For the forward modeling, we use a finite volume solver with a third-order weighted essentially non-oscillatory method (WENO3) for the cell-face fluxes and an implicit optimized second-order backward differentiation formulation (BDF2-opt) to handle the temporal evolution, combined with the URANS $k$-$\omega$ SST turbulence model. The efficient implementation of these methods enables accurate and fast approximation of turbulent thermally-coupled flows and the nonlinear phase change process. However, the small space and time scales required to describe the phase-change phenomenon [42], together with the need to systematically repeat these simulations—as needed for the inverse problem—calls for using Reduced-Order Modeling (ROM) strategies. With that, we can reliably decrease the dimensionality of the problem while still preserving its fundamental characteristics. Well-established ROM techniques include the reduced-basis method [43], proper orthogonal decompositions [44], randomized decompositions for large scale systems [45], modern hybrid approaches based on machine learning techniques [46], high order dynamic mode decompositions [47], non-linear approaches [48], among many others (see e.g. [49, 50] for a review on ROM strategies).
@@
As an inverse technique, we employ a regularized least-squares approach based on the so-called Parametrized Background Data Weak (PBDW) method, see, e.g., [35, 51, 52, 53]. In this approach, the optimization problem to be solved for the state estimation is based on the projection of the physics on a reduced-dimensional space that must reproduce the observed data in a weak form. Application of the PBDW in biomedical problems has been recently presented in [54, 55, 56, 57]. In this work, we use a methodology inspired by the PBDW to assimilate the temperature fields in the liquid-solid phase change, including turbulent flow conditions. In particular, we employ the ROM-based component of PBDW, which is equivalent to the full PBDW approach with infinite model regularization, as explained in [35], assuming no bias on the physical model. Additionally, we combine the state estimation with a Greedy algorithm for the optimal placement of temperature sensors. Greedy algorithms have been used for experimental design in the past years. In [58], a Bayesian design is used for seismic full waveform inversion. In [59], a greedy algorithm was employed using a minimal variance error for parameter estimation. In [60], optimal sensors are computed targeting directly a mean squared error between the model and measurements. Greedy methods have been combined with the PBDW in, e.g., [61, 62, 35]. A recent related work is [63], where the problem is posed so that an optimal stabilization of the inverse approach is acquired and applied in a Hamiltonian system for wave-type phenomena. In [64], sparse measurement locations are sought from a dictionary of available positions using a Wasserstein barycenter criteria. In contrast with these works, our proposal solely relies on available a priori results for the regularized least-squares reconstruction, prioritizing sensor locations that maximize the observability of a static reduced-order model.
@@
In summary, this article proposes a temperature estimation framework from optimally placed synthetic data and a regularized least-squares estimator through a physics-informed projection, using turbulent simulations of the airflow surrounding the freezing food. The temperature estimation algorithm is proven valid for extrapolating fields in non-observable regions, keeping good reconstruction errors due to a properly designed sensor placement strategy. As a further application of the fully reconstructed field, we also show the capabilities of the algorithm to predict temperature-related quantities, such as the local freezing curve and the local freezing rate, aspects of high interest in food engineering. This proof-of-concept article will specifically use synthetically generated measurements from a thermal camera.
@@ 2 Mathematical modeling of the conjugate food freezing process
In this work, we consider a two-dimensional (2D) setup for the forward simulations and the inverse reconstruction, as depicted in Fig. 1$(a)$. That geometrical simplification is based on the symmetry of the three-dimensional problem. We also performed a validation with experimental data previously published, which confirms that the use of the 2D model is adequate for the physical situation studied [65]. In that article, the importance of using ad-hoc turbulence models to describe the freezing evolution was discussed in detail, with URANS strategies being sufficiently accurate and efficient to reproduce experimental results. Since the key novelty of this article is the inverse estimation, we accept the 2D direct solution and use the 2D results as the ground truth values. The inverse methodology, without the inclusion of the greedy algorithm, was systematically tested on 3D scenarios with good outcomes in [56, 54]. In this concern, we do not restrict the analysis to the two dimensional case due to methodological limitations, since there are none, but rather because we would like to introduce a complete minimal reconstruction pipeline for a realistic scenario, still with a 2D set-up.
@@ 2.1 Model setting
Even though humidity and ice formation are important for freezing applications [66, 67, 68], including the mass transfer equation and more sophisticated phase-change models to characterize the airflow-food interaction is highly challenging from the forward modeling viewpoint. Also, including thermodynamic relations to compute the water condensation when the vapor concentration is higher than the saturated vapor concentration is necessary to upgrade the physical modeling of freezing applications. Since this work is focused on inverse modeling and temperature state estimation, with the finite volume solutions as input data, the mathematical model defined to represent the salmon freezing process only considers the momentum, continuity, and energy conservation laws. The particular treatment used to define the airflow and the food is defined separately in the following subsections.
We model the freezing process of food inside a convective freezer that circulates cold air between the cabinet and the refrigeration unit through inlet and outlet ducts. The conjugate dynamics of the process results from the interplay of forced heat convection in the turbulent airflow and the nonlinear heat transfer coupled with the liquid-to-solid phase change process within the food. This setup is illustrated in Fig. 1.
The freezer is assumed to have a cabinet of width $W=0.35$ m, height $H=0.4$ m, and depth $D=0.45$ m. Cold air coming from the refrigeration unit is forced into the cabinet by a fan at a constant velocity $U_{in}$ and temperature $T_{C}$ through an inlet duct of height $H_{in}=0.12$ m, and it leaves the cabinet through an outlet duct of height $H_{out}=0.06$ m. An idealized hexahedral food slab of width $W_{f}=0.08$ m, height $H_{f}=0.03$ m, and depth $D_{f}=0.08$ m is placed on a glass shelf of width $W_{s}=0.25$ m and thickness $d_{s}=0.002$ m, located at the mid-height of the cabinet. Also, we assume that the walls of the cabinet exchange heat with the environment by conduction through the insulated layers and by natural convection with the surrounding air. The rear wall that separates the cabinet from the refrigeration unit is considered adiabatic. In what follows, we will denote with $\Omega_{f}\subset\mathbb{R}^{2}$ and $\Omega_{s}\subset\mathbb{R}^{2}$ the fluid-flow and food domains, respectively, and $[0,t_{f}]$ is the physical time interval considered for the simulation.
Figure 1: (a) Sketch of the physical domain for the food freezing process with a brief description of the freezing phenomena. (b) Details of the two-dimensional section considered in this study.
@@ 2.2 Airflow dynamics
$\rho\partial_{t}T+\rho\boldsymbol{u}\cdot\nabla T-\nabla\cdot[(\lambda/C_{p}+\sigma_{T}\mu_{t})\nabla T]=0\qquad\text{in}\ \Omega_{f}\times(0,t_{f}].$
In Eq. (1) $\rho$, $\lambda$, $C_{p}$, and $\mu$ denote the density, thermal conductivity, specific heat, and viscosity of the air, respectively. These physical properties are assumed as constant at a reference temperature of $T_{ref}=0$$\degree$°C. This reference temperature was determined as the average between the maximum and minimum temperature values during the process. Then, Ideal Gas and Sutherland’s Laws were used to compute the values of density (1.2930 kg/m3), thermal conductivity (0.02436 W/mK1), and viscosity (1.3316 $\times$ $10^{-5}$ Ns/m2), while the specific heat was assumed as constant with a value of 1003.68 J/(kgK) and the thermal expansion coefficient was computed as $T_{ref}^{-1}$.
The source term in the momentum equation is defined as:
@@
and includes the buoyancy and the turbulent source terms, where $\beta_{ref}$ is the air’s thermal expansion coefficient and $k$ is the turbulent kinetic energy. The additional turbulent viscosity $\mu_{t}$ introduced by the URANS method is described by the $k$-$\omega$ SST model. The model attempts to predict turbulence and close the problem using two partial differential equations for two variables: $k$ and $\omega$. The first variable is the turbulence kinetic energy, while the second is the specific rate of dissipation. The turbulence model is based on two transport equations defined as:
$\rho\partial_{t}k+\rho\boldsymbol{u}\cdot\nabla k-\nabla\cdot(\mu+\sigma_{k}\mu_{t})\nabla k=\tilde{P}_{k}-\beta^{*}\rho k\omega\qquad\text{in}\ \Omega_{f}\times(0,t_{f}],$ (2)
$\rho\partial_{t}\omega+\rho\boldsymbol{u}\cdot\nabla\omega-\nabla\cdot(\mu+\sigma_{\omega}\mu_{t})\nabla\omega=\frac{\rho\gamma}{\mu_{t}}\tilde{P}_{k}-\rho\beta\omega^{2}+C_{k\omega}\qquad\text{in}\ \Omega_{f}\times(0,t_{f}].$
By using the two turbulence equations, the turbulent viscosity is defined as a function of the turbulent kinetic energy and specific turbulence dissipation rate ($\omega$) as:
$\mu_{t}=\rho\,\frac{a^{*}k}{\max(a_{1}\omega,SF_{2})}\,.$ (3)
@@ 2.3 Food freezing process
The internal convection of the liquid water contained in the muscle fibers and connective tissue of salmon can be neglected at the macro-level [70, 71, 72]. To account for the effect of latent heat during the phase change process in the salmon slab, the thermal properties (the density $\rho_{s}$, the specific heat capacity $C_{s}$, and the thermal conductivity $\lambda_{s}$) are considered as temperature-dependent, according to the effective heat capacity method [73, 74].
The heat transfer and the liquid-solid phase change process in the food slab are described by a non-linear energy equation, assuming an incompressible and homogeneous material neglecting mass transfer, such that:
$\partial_{t}(\rho_{s}C_{s}T_{s})-\nabla\cdot(\lambda_{s}\nabla T_{s})=0\qquad\text{in}\ \Omega_{s}\times(0,t_{f}].$ (4)
The temperature variability of the properties is depicted in Fig. 2, assuming third-order fitting with (piecewise-constant) temperature-dependent coefficients:
@@ 2.4 Boundary, initial, and interface conditions
The inverse pipeline of the present work requires a dataset of snapshots from a set of time-dependent direct simulations with a variable space of parameters. Therefore, the magnitudes of the main boundary conditions are varied within a fixed range: $U_{in}\in[0.15,0.25]$ m/s, $T_{C}\in[-18,-26]\ ^{\circ}\text{C}$, $T_{ext}\in[18,26]\ ^{\circ}\text{C}$, and $h_{ext}\in[0.4,1.2]\ \text{W/}\text{m}^{2}\text{K}$. The detailed procedure for constructing this five-dimensional set is explained in Section 5.
Since natural convection is important in the problem at hand, we need to define the Rayleigh number for each studied case:
$Ra=\frac{|\boldsymbol{g}|\beta(T_{C}-T_{ext})L_{c}^{3}}{\nu\alpha},$
where for each case $T_{C}$ and $T_{ext}$ change and the thermophysical properties must be redefined using $T_{C}$. The characteristic length is defined as the height of the freezer cabinet in all cases. Using this Rayleigh definition, we can confirm that the values are very high ($Ra>10^{10}$), which is the reason for including a turbulence model in the direct simulations.